If the edge of a cube increased at the rate of 60cm per second at what rate the volume is increasing when the edge is 90cm

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi,


ur queswtion is a simple use of derivatives

 

rate of increase of side say 'a', is da/dt = 60


since volume of a cube  , v = a^3

so dv/dt = 3*a2 * da/dt

here dv/dt = rate of increase of volume


so dv/dt = 3* 90*90* 60

            = 1458000 cubic centimeter / sec